# Mellin‑Space Prony Representability of Linear Viscoelastic Models

### v1

@@ Abstract
Linear viscoelasticity is universally described by relaxation functions, which are continuous in relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations. While Laplace-domain rational moduli admit finite Prony representations, using Mellin space analysis requires a deeper structural criterion. This work proves that finite Prony series representation holds if and only if the arithmetic pole lattices of Gamma factors in the Mellin transform of the complex modulus align exactly with those of a trial kernel, with residues satisfying decoupled first-order recurrences along aligned sublattices. This maximal criterion classifies classical models (Maxwell, standard linear solid) as finitely representable and fractional models (power-law, Cole–Cole, Zener, Gaussian) as transcendentally representable via infinite Prony ladders, yielding a complete analytical pole-lattice taxonomy of viscoelastic material models.
@@ 1 Introduction
A material is linear viscoelastic if stress is a linear time-dependent functional of strain. Linear viscoelasticity is universally described by relaxation functions, which are continuous distributions of relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations.
@@
With the caveat that “all models are wrong but some are useful” [1] – a stance universally adopted in statistics and data science —this works explores the utility of Prony series in describing linear viscoelasticty from complex-analytical perspective. The Prony series provides an approximate representation of any viscoelastic model through finite sums of exponential terms in its time-dependent relaxation moduli. A comprehensive recent review of Prony-series applications in viscoelasticity is given in [2].
@@
From a dynamical-systems perspective, finding a linear viscoelastic representation is equivalent to constructing a mechanical network—a finite arrangement of springs and dashpots—whose input-output behavior reproduces the prescribed complex=valued dynamic modulus $G^{*}(\omega)$. Such networks generate relaxation spectra that are finite sums of Dirac masses (finite Prony series) with rational transfer functions in the frequency (Fourier/Laplace) variable. The inverse question—whether a given modulus admits a finite network representation—therefore translates into a classical realization problem: does the system possess a finite-dimensional state-space description, or equivalently, does its transfer function exhibit a pole structure consisting of a finite set of simple poles on the negative real axis? The Mellin-transform framework developed here recasts this condition as an algebraic constraint on the pole lattices of the transformed kernel, thereby linking the geometry of the complex-plane singularities directly to the existence of finite spring–dashpot analogous networks.
@@
Classical spring-dashpot networks (Maxwell, Kelvin-Voigt, Standard Linear Solid, and their finite generalizations) yield linear ordinary differential equations for the constitutive relations. Furthermore, their complex moduli $G^{*}(\omega)$ are rational functions of the imaginary spectral variable $i\omega$ in the Fourier/Laplace domains, admitting exact Prony representations with finitely many exponential terms. That is, in the frequency domain their spectrum contains a finite number of poles. Fractional constitutive laws, by contrast, involve fractional memory-dependent, non-local (e.g. containing combinations of differential and integral operators, i.e. ”differintergrals”) constitutive laws [3]. Many soft materials exhibit power-law viscoelastic responses across typical observation bandwidths, modeled by continuous distributions of relaxation times that reflect their fractal internal structure [4]). Structurally, power-law models, Cole-Cole, and Havriliak-Negami kernels generate non-rational dynamic moduli $G^{*}(\omega)$ with branch-cuts, implying infinite relaxation times distributions.
@@
Recent works surveys Prony-series approximations of fractional materials’ data  [2, 5]. Yet, these approximations remain ad hoc constructions for a given model. To understand the dichotomy between finite Prony models and non-local fractional models one needs to resolve the inverse problem: given a modeled $G^{*}(\omega)$, does it admit a faithful finite Prony representation? This question is numerically ill-posed: recovering $\{g_{k},\tau_{k}\}$ from frequency data suffers Hadamard instability, where noise amplifies into spurious oscillations, obscuring the discrete-vs-continuous distinction [6, 7]. Therefore, to establish the answer one needs to employ analytical tools.
@@ 2.1 Mellin transforms
    For a pure power-law $f(\tau)\sim\tau^{\beta}$, the Mellin transform is defined via analytic continuation from the strip $(0,1)\to[1,\infty)$
@@ 3.2 Mellin transform of the spectral kernel
$\tilde{K}(s)=\mathcal{M}_{\omega}\{K(\omega\tau)\}(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s),\qquad 0<\Re s<1,$ (8)
@@ 3.3 Constitutive equation in Mellin space
To avoid dealing with the Dirac $\delta$ distribution let $\Delta G^{*}(\omega)=G^{*}(\omega)-G_{\infty}$, the viscoelastic part of the complex modulus, and
@@ 4.1 Degenerate case
$\tilde{H}(s)=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s},$
where $c_{k}=g_{k}>0$ and $\lambda_{k}=\ln\tau_{k}\in\mathbb{R}$.
@@
($\Leftarrow$) Suppose $\tilde{H}(s)=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}\in\mathbb{R}$ and $c_{k}\neq 0$. The inverse Mellin transform along a vertical line $\Re s=\sigma$ in the strip of convergence yields
$H(\tau)=\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}\tilde{H}(s)\,ds=\sum_{k=1}^{N}c_{k}\cdot\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds.$
The inner integral is the inverse Mellin transform of $e^{\lambda_{k}s}$. Since $e^{\lambda_{k}s}=\mathcal{M}\{\delta(\ln\tau+\lambda_{k})\}(s)$ in the distributional sense, we obtain formally
$\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds=\delta(\ln\tau+\lambda_{k})=\tau_{k}\,\delta(\tau-\tau_{k}),$
@@
where $\tau_{k}=e^{-\lambda_{k}}$. Hence $H(\tau)=\sum_{k}c_{k}\tau_{k}\delta(\tau-\tau_{k})$, which is a finite Prony series with $g_{k}=c_{k}\tau_{k}$. ∎
@@ 4.2 Generic case
$G(t)=\int_{0}^{\infty}a(t/\tau)\,\tau^{-(\beta+1)/\alpha}\,\frac{d\tau}{\tau},$
between a finite Prony series $a(t/\tau)$ and the power-law kernel $\tau^{-(\beta+1)/\alpha}$ whose Mellin transform is
$\mathcal{M}\{\tau^{-(\beta+1)/\alpha}\}(s)=\frac{1}{\alpha}\Gamma(\alpha s+\beta).$
Thus the trial state factorization $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ emerges from a multiplicative convolution in the relaxation timescale-space $\tau$. Finite Prony exactness then reduces to the existence of (i) a finite Prony sum $a(t/\tau)$ whose Mellin transform recovers the prescribed modulation $A(s)$ and (ii) the existence of power-law whose subordination yields $G(t)$ on the analyticity strip. This combination amounts to the fulfillment of the two technical conditions identified in Th. 1. Such an interpretation provides the analytic foundation for the ontological reality of finite-mode relaxation dynamics, emerging from the continuous power-law spectrum, through subordination in relaxation timescale space. Mutatis mutandis, if $a(t/\tau)$ is instead allowed to be an infinite Prony series
$a(u)=\sum_{k=1}^{\infty}c_{k}e^{-u/\tau_{k}},$
@@
When these conditions hold, the solution $\widetilde{H}(s)\in\mathcal{P}$ is unique up to a scale factor.
@@
    $\operatorname{Res}_{s=n}\tilde{K}(s)=\pi ie^{-i\pi n/2}\cdot(-1)^{n}=\pi i(-1)^{n}e^{-i\pi n/2}.$
@@
$K(s)\tilde{H}(-s)=E(s)-\sum_{j=1}^{J}\sum_{k=0}^{\infty}\frac{R_{j,k}K(-\lambda_{j}^{(k)})}{s+\lambda_{j}^{(k)}}+\sum_{n\in\mathbb{Z}}\frac{\pi i(-1)^{n}e^{-i\pi n/2}\tilde{H}(-n)}{s-n},$ (12)
@@
Now, $K(s_{n})=\frac{\pi ie^{-i\pi s_{n}/2}}{\sin(\pi s_{n})}$ and $\rho_{n}=A(s_{n})(-1)^{n}/(\alpha n!)$. Therefore
@@
    $\int_{0}^{\infty}H(\tau)\frac{d\tau}{t}=G_{0}-G_{\infty}.$
@@ 5 Transcendental Prony Representation
Suppose that a given $G(t)$ is completely monotone, but its complex modulus $G^{*}(\omega)$ does not admit a finite Prony representation, that is no $\tilde{H}(s)\in\mathcal{P}$ satisfies the constitutive equation $\tilde{G}(s)=K(s)\tilde{H}(-s)$. Then $G(t)$ admits an exact representation by the infinite Prony ladder:
$H(\tau)=\sum_{k=-\infty}^{\infty}g_{k}\tau_{k}\delta(\tau-\tau_{k}),\qquad\tau_{k}=\tau_{0}q^{k},\;q>1,$
where the weights $g_{k}$ are determined by the continuous relaxation spectrum $H_{c}(\tau)$ existing by Bernstein’s theorem, and the corresponding generalized Mellin kernel satisfies the constitutive equation distributionally. Such representation is called transcendental.
@@
Define the discrete approximations
@@
For the Mellin transform, standard quadrature theory on a logarithmic grid gives
$\lim_{N\to\infty}\mathcal{M}\{H_{N}\}(s)=\mathcal{M}\{H_{c}\}(s)=\tilde{H}_{c}(s)$
for $s$ in the common strip of convergence. Since $\tilde{G}(s)=K(s)\tilde{H}_{c}(-s)$ holds for the continuous spectrum, and $\mathcal{M}\{H_{N}\}(s)\to\tilde{H}_{c}(s)$ pointwise, the discrete spectra $H_{N}$ converge distributionally to a representation of $H_{c}(\tau)$ as an infinite sum of Dirac masses—the infinite Prony ladder. In the limit, this ladder satisfies the constitutive equation in the sense of distributions. ∎
Corollary 1 (Entire non-exponential spectra are transcendental).
@@
Suppose $\exp(P(s))$ were a finite exponential sum: $\exp(P(s))=\sum_{k=1}^{N}c_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}$. Taking the logarithmic derivative gives
$P^{\prime}(s)=\frac{\sum_{k}c_{k}\lambda_{k}e^{\lambda_{k}s}}{\sum_{k}c_{k}e^{\lambda_{k}s}}.$
@@ 5.1 Matching of arithmetic progressions
Lemma 1 (Mismatch of arithmetic progressions).
@@
    If $\alpha=p/q$ with $p,q\in\mathbb{N}$, $\gcd(p,q)=1$ and $p\neq q$, then $\Lambda_{\alpha,\beta}\cap\mathbb{Z}$ is either empty or an arithmetic progression with spacing p/|pq|p/|p-q|, but in any case $\Lambda_{\alpha,\beta}\not\subseteq\mathbb{Z}$.
@@
    If $\alpha=1$ (the integer‑spacing case), then $\Lambda_{1,\beta}\subseteq\mathbb{Z}$ if and only if $\beta\in\mathbb{Z}$.
@@
If $\alpha$ is irrational, this linear Diophantine equation in $m,k$ can have at most one solution.
If $\alpha=p/q$ rational, the equation becomes $qm+pk=-q\beta$. This is a linear Diophantine equation in $m,k$. Its solutions, if any, form an arithmetic progression with step $\operatorname{lcm}(q,p)/q=p/\gcd(p,q)=p$ in $m$. Thus the integers $s_{m}$ that appear are spaced by $p$, not 1, so the progression cannot cover all integers unless $p=1$.
If $\alpha=1$, then $s_{m}=-\beta-m$. These are all integers precisely when $\beta$ is an integer.
The final statement follows from the previous cases: only when $\alpha=1$ can the progression have spacing 1, and only when $\beta\in\mathbb{Z}$ does it lie exactly on the integers. ∎

### v2

@@ Abstract
Linear viscoelasticity is universally described by relaxation functions, which are continuous in relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations. While Laplace-domain rational moduli admit finite Prony representations, using Mellin space analysis requires a deeper structural criterion. This work proves that finite Prony series representation holds if and only if the arithmetic pole lattices of Gamma factors in the Mellin transform of the complex modulus align exactly with those of a trial kernel, with residues satisfying decoupled first-order recurrences along aligned sublattices. This maximal criterion classifies classical models (Maxwell, standard linear solid) as finitely representable and fractional models (power-law, Cole–Cole, Zener) and the log-normal one as transcendentally representable via infinite Prony ladders, yielding a complete analytical pole-lattice taxonomy of viscoelastic material models.
@@ 1 Introduction
A material is linear viscoelastic if stress is a linear time-dependent functional of material strain. Linear viscoelasticity is universally described by relaxation functions, which are continuous distributions of relaxation times. Yet experimentally, measurements are necessarily discrete and band-limited, creating a fundamental gap between the continuous mathematical description and empirical observations.
@@
With the caveat that “all models are wrong but some are useful” [1] – a stance universally adopted in statistics and data science —this works explores the utility of Prony series in describing linear viscoelasticty from complex-analytical perspective. The Prony series provides an approximate representation of any viscoelastic model through finite sums of exponential terms in its time-dependent relaxation shear moduli. A comprehensive recent review of Prony-series applications in viscoelasticity is given in [2].
@@
From a dynamical-systems perspective, finding a linear viscoelastic representation is equivalent to constructing a mechanical network–a finite arrangement of springs and dashpots–whose input-output behavior reproduces the prescribed complex-valued dynamic modulus $G^{*}(\omega)$. Such networks generate relaxation spectra that are finite sums of Dirac masses (finite Prony series) with rational transfer functions in the frequency (Fourier/Laplace) variable. The inverse question—whether a given modulus admits a finite network representation—therefore translates into a classical realization problem: does the system possess a finite-dimensional state-space description, or equivalently, does its transfer function exhibit a pole structure consisting of a finite set of simple poles on the negative real axis? The Mellin-transform framework developed here recasts this condition as an algebraic constraint on the pole lattices of the transformed kernel, thereby linking the geometry of the complex-plane singularities directly to the existence of finite spring–dashpot analogous networks.
@@
Classical spring-dashpot networks (Maxwell, Kelvin-Voigt, Standard Linear Solid, and their finite generalizations) yield linear ordinary differential equations for the constitutive relations. Furthermore, their complex moduli $G^{*}(\omega)$ are rational functions of the imaginary spectral variable $i\omega$ in the Fourier/Laplace domains, admitting exact Prony representations with a finite number of exponential terms. That is, in the frequency domain their spectrum contains a finite number of poles. Fractional constitutive laws, by contrast, involve fractional memory-dependent, non-local (e.g. containing combinations of differential and integral operators, i.e. ”differintergrals”) constitutive laws [3]. Many soft materials exhibit power-law viscoelastic responses across typical observation bandwidths, modeled by continuous distributions of relaxation times that reflect their fractal internal structure [4]). Structurally, power-law models, Cole-Cole, and Havriliak-Negami kernels generate non-rational dynamic moduli $G^{*}(\omega)$ with branch-cuts, implying infinite relaxation times distributions.
@@
Recent works survey Prony-series approximations of fractional materials’ data  [2, 5]. Yet, these approximations remain ad hoc constructions for a given model. To understand the dichotomy between finite Prony models and non-local fractional models one needs to resolve the inverse problem: given a modeled $G^{*}(\omega)$, does it admit a faithful finite Prony representation? This question is numerically ill-posed: recovering $\{g_{k},\tau_{k}\}$ from frequency data suffers Hadamard instability, where noise amplifies into spurious oscillations, obscuring the discrete-vs-continuous distinction [6, 7]. Therefore, to establish the answer one needs to employ analytical tools.
@@ 2.1 Mellin transforms
    For a pure power-law $f(\tau)\sim\tau^{\beta}$, the Mellin transform is defined via analytic continuation from the strip $(0,1)\to[1,\infty)$ of the range-limited integral
@@ 2.2 The Finite Prony Class $\mathcal{P}$
When introducing the above definition some remarks are in order.
The class $\mathcal{P}$ contains both finite discrete Prony spectra (exponential sums) and continuous spectra representable via Gamma ratios. Throughout this work, the question of finite Prony representability refers specifically to the former: whether a model can be expressed as a finite sum of exponentials/atomic measures, i.e., belongs to the subclass characterized by Proposition 1.
@@ 3.2 Mellin transform of the spectral kernel
$\tilde{K}(s)=\mathcal{M}_{\omega}\{K(\omega\tau)\}(s)=-\pi e^{i\pi(1-s)/2}\csc(\pi s),\qquad-1<\Re s<0,$ (8)
@@ 3.3 Constitutive equation in Mellin space
To avoid dealing with the Dirac $\delta$ distribution let $\Delta G^{*}(\omega)=G^{*}(\omega)-G_{\infty}$, be the viscoelastic part of the complex modulus, and
@@ 4.1 Degenerate case
$\tilde{H}(s)=\sum_{k=1}^{N}g_{k}e^{\lambda_{k}s},$
where $\lambda_{k}=\ln\tau_{k}\in\mathbb{R}$ .
@@
($\Leftarrow$) Suppose $\tilde{H}(s)=\sum_{k=1}^{N}g_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}\in\mathbb{R}$ and $g_{k}>0$. The inverse Mellin transform along a vertical line $\Re s=\sigma$ in the strip of convergence yields
$H(\tau)=\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}\tilde{H}(s)\,ds=\sum_{k=1}^{N}g_{k}\cdot\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds.$
The inner integral is the inverse Mellin transform of $e^{\lambda_{k}s}$. Since $e^{\lambda_{k}s}=\mathcal{M}\{\delta(\ln\tau-\lambda_{k})\}(s)$ in the distributional sense, we obtain formally
$\frac{1}{2\pi i}\int_{\sigma-i\infty}^{\sigma+i\infty}\tau^{-s}e^{\lambda_{k}s}\,ds=\delta(\ln\tau-\lambda_{k})=\tau\,\delta(\tau-\tau_{k}),,$
@@
where $\tau_{k}=e^{\lambda_{k}}$. Hence $H(\tau)=\sum_{k}g_{k}\tau_{k}\delta(\tau-\tau_{k})$, which is a finite Prony series. ∎
@@ 4.2 Generic case
$G(t)=\int_{0}^{\infty}a(t/\tau)\,\tau^{\beta/\alpha}e^{-\tau^{1/\alpha}}\frac{d\tau}{\tau},$
between a finite Prony series $a(t/\tau)$ and the trial state gamma kernel $\tau^{\beta/\alpha}e^{-\tau^{1/\alpha}}$ whose Mellin transform is
$\mathcal{M}\{\tau^{\beta/\alpha}e^{-\tau^{1/\alpha}}\}(s)={\alpha}\Gamma(\alpha s+\beta).$
Thus the trial state factorization $\tilde{G}(s)=A(s)\Gamma(\alpha s+\beta)$ emerges from a multiplicative convolution in the relaxation temporal scale-space $\tau$. Finite Prony exactness then reduces to the existence of (i) a finite Prony sum $a(t/\tau)$ whose Mellin transform recovers the prescribed modulation $A(s)$ and (ii) the existence of power-law whose subordination yields $G(t)$ on the analyticity strip. This combination amounts to the fulfillment of the two technical conditions identified in Th. 1. Such an interpretation provides the analytic foundation for the ontological reality of finite-mode relaxation dynamics, emerging from the continuous power-law spectrum, through subordination in relaxation timescale space. Mutatis mutandis, if $a(t/\tau)$ is instead allowed to be an infinite Prony series
$a(u)=\sum_{k=1}^{\infty}g_{k}e^{-u/\tau_{k}},$
@@
When these conditions hold, a solution $\widetilde{H}(s)\in\mathcal{P}$ exists. If, in addition, the entire factor $Q(s)=\exp(P(s))$ in the representation $\widetilde{H}(s)=H_{\Gamma}(s)Q(s)$ is required to be of exponential type with
$\limsup_{y\to\pm\infty}\frac{|\log|Q(iy)||}{|y|}<\pi,$
then the solution is unique up to an overall multiplicative constant (scale factor), which is fixed by the physical balance condition
$\int_{0}^{\infty}H(\tau)\frac{d\tau}{\tau}=G_{0}-G_{\infty}.$
@@
    $\operatorname{Res}_{s=n}\tilde{K}(s)=-e^{-i\pi n/2}\cdot(-1)^{n}=-(-1)^{n}e^{-i\pi n/2}.$
@@
$K(s)\tilde{H}(-s)=E(s)-\sum_{j=1}^{J}\sum_{k=0}^{\infty}\frac{R_{j,k}K(-\lambda_{j}^{(k)})}{s+\lambda_{j}^{(k)}}+\sum_{n\in\mathbb{Z}}\frac{-(-1)^{n}e^{-i\pi n/2}\tilde{H}(-n)}{s-n},$ (12)
@@
Now, $K(s_{n})=\frac{-e^{-i\pi s_{n}/2}}{\sin(\pi s_{n})}$ and $\rho_{n}=A(s_{n})(-1)^{n}/(\alpha n!)$. Therefore
@@
    $\int_{0}^{\infty}H(\tau)\frac{d\tau}{\tau}=G_{0}-G_{\infty}.$
@@ 5 Transcendental Prony Representation
Suppose that $G(t)$ is completely monotone, but its relaxation spectrum does not admit a finite Prony representation. Let $H_{c}(\tau)$ be the continuous relaxation spectrum whose existence is guaranteed by Bernstein’s theorem, so that
Choose a geometric progression $\tau_{k}=\tau_{0}q^{k}$ with $q>1$ and set $g_{k}=H_{c}(\tau_{k})\ln q$. Then the discrete atomic measures
$d\mu_{N}:=\sum_{k=-N}^{N}g_{k}\tau_{k}\delta(\tau-\tau_{k})$
converge weakly to the continuous measure $d\mu=H_{c}(\tau)\frac{d\tau}{\tau}$ as $N\to\infty$; i.e., for every continuous bounded function $\varphi\in C_{b}(0,\infty)$,
$\lim_{N\to\infty}\int_{0}^{\infty}\varphi(\tau)\,d\mu_{N}=\int_{0}^{\infty}\varphi(\tau)H_{c}(\tau)\frac{d\tau}{\tau}.$
The infinite distributional sum
$H(\tau):=\sum_{k=-\infty}^{\infty}g_{k}\tau_{k}\delta(\tau-\tau_{k})$
is called a transcendental Prony ladder; it represents the continuous spectrum $H_{c}(\tau)$ in the sense of weak convergence of measures.
@@
Define the discrete measures
@@
For functions in the Mellin–Hardy space $\mathcal{H}^{p}(\mathbb{R}^{+})$ (or more generally, whenever the Mellin–Poisson summation formula holds exactly [13]), the Mellin transform satisfies the exact quadrature identity
$\lim_{N\to\infty}\mathcal{M}\{H_{N}\}(s)=\mathcal{M}\{H_{c}\}(s)=\tilde{H}_{c}(s),$
pointwise in the strip of convergence. Since $\tilde{G}(s)=K(s)\tilde{H}_{c}(-s)$ holds for the continuous spectrum, and $\mathcal{M}\{H_{N}\}(s)\to\tilde{H}_{c}(s)$, the infinite ladder $H(\tau)=\sum_{k=-\infty}^{\infty}g_{k}\tau_{k}\delta(\tau-\tau_{k})$ satisfies the constitutive equation distributionally. This yields an exact transcendental representation. ∎
Remark 3 (On the nature of convergence).
The convergence established in Theorem 2 is distributional (i.e weak convergence in the sense of measures). This means that for any continuous test function $\phi(\tau)$ with compact support on $(0,\infty)$,
$\lim_{N\to\infty}\int_{0}^{\infty}\phi(\tau)H_{N}(\tau)d\tau=\int_{0}^{\infty}\phi(\tau)H_{c}(\tau)\frac{d\tau}{\tau}.$
The limit object is the absolutely continuous measure $H_{c}(\tau)d\tau/\tau$, not a discrete sum of Dirac masses. The infinite Prony ladder therefore reproduces the continuous spectrum when integrated against suitable test functions, and consequently yields the correct complex modulus $G^{*}(\omega)$ for all $\omega>0$ via the constitutive equation.
This is the appropriate notion of convergence for the problem at hand: physical observables (stress, strain, complex modulus) are always measured through such test functions (e.g., finite frequency bands, finite time windows). In practice, any numerical implementation truncates the infinite ladder to a finite sum $H_{N}$ with sufficiently large $N$, and the distributional convergence guarantees that the error in any physically measurable quantity can be made arbitrarily small.
Corollary 1 (Entire, non-exponential spectra are transcendental).
@@
Suppose $\exp(P(s))$ were a finite exponential sum: $\exp(P(s))=\sum_{k=1}^{N}g_{k}e^{\lambda_{k}s}$ with distinct $\lambda_{k}$. Taking the logarithmic derivative gives
$P^{\prime}(s)=\frac{\sum_{k}g_{k}\lambda_{k}e^{\lambda_{k}s}}{\sum_{k}g_{k}e^{\lambda_{k}s}}.$
@@ 5.1 Matching of arithmetic progressions
Lemma 1 (Intersection of arithmetic progressions with integers).
@@
    If $\alpha=p/q$ with $p,q\in\mathbb{N}$, $\gcd(p,q)=1$ and $p\neq q$, then either $\Lambda_{\alpha,\beta}\cap\mathbb{Z}=\varnothing$, or it is an arithmetic progression with spacing $q$ in the $s$-variable.
@@
    If $\alpha=1$, then $\Lambda_{1,\beta}\subseteq\mathbb{Z}$ if and only if $\beta\in\mathbb{Z}$.
@@
If $\alpha$ is irrational, this linear Diophantine equation in $m,k$ can have at most one solution, proving (1).
If $\alpha=p/q$ with $p,q\in\mathbb{N}$, $\gcd(p,q)=1$, the equation becomes
$m+\frac{p}{q}k=-\beta\quad\Longleftrightarrow\quad qm+pk=-q\beta.$
This is a linear Diophantine equation in $m,k$. If a particular solution $(m_{0},k_{0})$ exists, the general solution is
$m=m_{0}+pt,\quad k=k_{0}-qt,\quad t\in\mathbb{Z}.$
Thus the indices $m$ yielding integer $s_{m}$ differ by $p$. The corresponding $s$-values are
$s=k=k_{0}-qt,$
which form an arithmetic progression with step $q$ in the $s$-variable. This proves (2).
If $\alpha=1$, then $s_{m}=-\beta-m$. These are all integers precisely when $\beta\in\mathbb{Z}$, proving (3).
Statement (4) follows from (2) and (3): when $\alpha\neq 1$, the step in $s$ is $q>1$, so the progression cannot be contained in $\mathbb{Z}$; when $\alpha=1$, containment requires $\beta\in\mathbb{Z}$. ∎